Circular No. 6546

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NEW X-RAY BURSTER T. Boller, F. Haberl, and W. Voges, Max-Planck-Institut für Extrater-

restrische Physik, Garching; L. Piro, Istituto Astrofisica Spaziale, CNR, Frascati; and J. Heise, Space Research Organization of the Netherlands, report: “Within the 10′ error box centered on the BeppoSAX position of the x-ray burst from Jan. 13.295011 UT (IAUC 6538), we have found the single x-ray source 1RXS J124938.0–590525, contained in the ROSAT All-Sky Survey Bright Source Catalogue (IAUC 6420) at the position α = 12^{h}49^{m}38^{s}, δ = −59^{o}05′.4 (equinox 2000.0; estimated 1σ error 9′′), which is 5′.6 north of the BeppoSAX position, with a countrate of 4.09 ± 0.27 counts/s (0.1−2.4 keV) observed during 1991 Aug. 4-5. During four ROSAT PSPC observations centered on α = 12^{h}47^{m}43^{s}, δ = −59^{o}41′.4, and totalling 12 400 s between 1993 July 11 and Aug. 27, two sources were detected within the BeppoSax error circle with the following positions, mean countrates and fluxes (assuming a power-law model): source 1, α = 12^{h}49^{m}39^{s}, δ = −59^{o}05′.2, 2.78 ± 0.02 counts/s (0.1–2.4 keV), or 1.7 × 10^{−10} erg cm^{−2} s^{−1} [for a power-law-fit value of the hydrogen column density N_{H} = (2.9 ± 0.9) × 10^{21} cm^{−2} and photon index 2.1 ± 0.2]; source 2, α = 12^{h}49^{m}27^{s}, δ = −59^{o}13′.3 (3′.9 south of the BeppoSAX position), 0.08 ± 0.002 counts/s (0.1–2.4 keV) or 2.2 × 10^{−12} erg cm^{−2} s^{−1} [for power-law-fit N_{H} = (0.4 ± 0.1) × 10^{21} cm^{−2} and photon index 2.4 ± 0.3]. Source 2 is probably a foreground object because of the relatively-low-power-law-fit N_{H} value. Source 1 shows variability by ± 20 percent during the pointing. Source 2 is not detected in the ROSAT all-sky survey. The upper limit value of 0.03 counts/s implies a change by more than a factor of 2 between the survey and pointing observations. The ROSAT source 1 (1RXS J124938.0–590525), as well as the BeppoSAX burster, are very likely associated with 4U1246−58. Flaring activity from this source was seen in the past — but, to our knowledge, a strong and short burst such as that observed by BeppoSAX (IAUC 6538) was never reported. The flat-top light curve observed in this event suggests a type-1 Eddington-limited burst.”

SUPERNOVA 1997B IN IC 438 Corrigendum. M. Cavagna and P. Sicoli report that their measurement

on IAUC 6537 was incorrect. The proper position is refined to read α = 5^{h}53^{m}02^{s}.97, δ = −17^{o}52′23′′.5.

1997 January 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green